OpenStudy (anonymous):

how to find f(1.5) from a calculator? HELP!

OpenStudy (anonymous):

You need to know what the function f represents

OpenStudy (anonymous):

what do you mean? that's exactly what I am asking. I mean what the symbol in the calculator?

OpenStudy (anonymous):

There is no f symbol in the calculator :D

OpenStudy (anonymous):

then how do we calculate it? everyone in my class was doing so. :/

OpenStudy (anonymous):

Maybe is that in hexadecimal? need to switch calculator to hex mode

OpenStudy (anonymous):

I have a fx-991MS calculator. There's no such mode.

OpenStudy (anonymous):

Then that will make it harder.....especially with the 1.5. Can't even find a calculator that will do heximal? points... could convert from hex back to decimal the f is easy that is a 15 one is just one...but fractions I don't see much on the web on fractions but it might mean .8??? so in decimal maybe and I do mean maybe it means 15 * 1.8 = 27 convert back to hex and get 1B???? But that is pretty sketchy. I tried 2 calculators here at home and neither of them will do fractional hex. And if your class is doing hex/octa/binary conversions a calculator with those modes is gonna be kind of important.

OpenStudy (anonymous):

hmm I don't know why you're taking the hexadecimal? I don't think its relevant. I think you didn't understand my question. I need to find the function of (1.5). That's it. I am pretty sure most of the students the sa,e calculator as mine. So there's another way out I suppose.

OpenStudy (anonymous):

except f(1.5) is not a function. It is part of a function but there has to be more after the f(x) part

OpenStudy (anonymous):

Well the question is: \[f(x)=x-1/x ^{2}+1\]

OpenStudy (anonymous):

So in order to find f(x) we input x= 1.5,1.1,1.01,1.001 etc etc.

OpenStudy (anonymous):

hehe ok then to solve the function for f(1.5) replace each of the x's with 1.5 (1.5)-1/(1.5)^2 +1 .....

OpenStudy (anonymous):

that looks like you are using a table to find a limit?

OpenStudy (anonymous):

LOL Ohhh okayyyy. Yeah I guess so. Our prof. called it the calculator method lol

OpenStudy (anonymous):

Thanks, buddy.

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

so ya just enter the different values into the equation wherever there is an x. I'll do a simpler function x+2 but in the values for x and write down the answer 1.5 + 2 =3.5 1.1 + 2 =3.1 1.01 + 2 =3.01 and you will see the answers get closer and closer to some value in this case 3 so the limit as x -> 1 f(x)=x+2 is 3

OpenStudy (anonymous):

and one last thing if you are still looking at this you will have three parts to this function the limit that x is approaching, the f(x) part and then the equation part

OpenStudy (anonymous):

hey if you're there, do you know what we will do if lim x approaches to \[\Pi\]